-6x+6y=12 and 6x+y=-12
Answer by Maths68(1474)

-6x+6y=12.........(1)
6x+y=-12...........(2)
Add (1) and (2)
-6x+6y=12
6x+y=-12
-----------
7y=0
7y/7=0/7
y=0
Put the value of y in (2)
6x+y=-12
6x+0=-12
6x=-12
6x/6=-12/6
x=-2
Solution set (-2, 0)
Check
Put the values of x and y in any of two equations (1) or (2)
-6x+6y=12.........(1)
-6(-2)+6(0)=12.........(1)
12+0=12
12=12
6x+y=-12...........(2)
6(-2)+0=-12
-12=-12